a 12ft board is cut into 2 pieces so that one piece is 8 feet longer than 3 times the shorter piece. If the shorter piece is x feet long, find the lengths of both pieces
Answer by Maths68(1474)

Shorter piece = x feet
Longer piece = 12-x feet
Longer piece=3 times (shorter piece) + 8
12-x=3x+8
-x-3x-8-12
-4x=-4
-4x/-4=-4/-4
x=1
Shorter piece = x = 1 foot
Longer piece = 12-x = 12-1= 11 feet